SPECTROSCOPY, DISSOCIATION DYNAMICS AND POTENTIAL ENERGY SURFACES FOR CN(A)-Ar

JIANDE HAN and \underline{MICHAEL C. HEAVEN}, Department of Chemistry, Emory University, Atlanta, GA 30322; UDO SCHNUPF, Department of Agriculture, ARS/NCAUR/PPL, Peoria, IL 61604.

The A^{2}Π-X^{2}Σ^{+} band system of CN-Ar has been examined using fluorescence depletion and action spectroscopy techniques. Eight vibronic bands of the complex were observed in association with the monomer 3-0 transition. Pump-probe measurements were used to characterize CN(A^{2}Π_{3/2}, v=3) fragments from direct photodissociation of CN(A^{2}Π, v=3)-Ar, and CN(X^{2}Σ^{+}, v=7) fragments from CN(A^{2}Π, v=3)-Ar predissociation. The latter showed a marked preference for population of positive parity diatomic rotational levels. Bound state calculations were used to assign the A-X bands, and to obtain fitted potential energy surfaces for the A state. The average potential obtained from fitting had a well depth of D_{e}=137.8 cm^{−1}. High-level ab initio calculations were used to obtain equilibrium Jacobi coordinates of θ_{e} = 94^{∘} and R_{e}=7.25 bohr. The near symmetric character of the fitted potential energy surface was consistent with the symmetry preference observed in the predissociation dynamics.